OpenStudy (anonymous):

It takes pipe A 20 days to fill a fish pone. Pipe B takes 15 days find how long it takes both pipes together to fill the pond.? How do you set up the problem, that's what i need help with.

OpenStudy (anonymous):

................sigh. A needs 20 days B needs 15 days per day A can do 1/20 job. per day B can do 1/15 of job. days(1/20) + days(1/15) = 1

OpenStudy (anonymous):

1/20+1/15 3/60+4/60 = 7/60

OpenStudy (anonymous):

a=20a=v b=15b=v da+db=v

OpenStudy (anonymous):

royal..7/60 is the part of pone fill by both A & B in a day. so the totl no of days taken = 60/7=8 days 13 hrs.

OpenStudy (anonymous):

Yes, you are right. I missed that part.

OpenStudy (anonymous):

thanx royal